OpenStudy (anonymous):

School day hrs of 4hrs, 6.5hrs, 3.5hrs, and 6.5. How long must friday hrs be in order to average at least 5.5 hrs for the week.

Parth (parthkohli):

Mean{4,6.5,3.5,6.5+x} = (20.5 + x)/5 = 5.5 => 20.5 + x = 27.5 => x = 7

Parth (parthkohli):

At least 7 hours

OpenStudy (anonymous):

no of working days = 5 total no of working hours if average is 5.5hrs\[=5.5\times5=27.5 hrs\] total no of hours engaged on \[4 + 6.5 + 3.5 + 6.5 + Friday = 20.5 + friday\] Hours engaged on friday = \[27.5 - 20.5 = 7 hours\]